OpenStudy (anonymous):

How do u solve y=(x-3)^2-4 ???? Someoneee help meeee pleaseeee

OpenStudy (cwrw238):

do you want to find x in terms of y?

OpenStudy (anonymous):

Yes. I keep getting y=x^2-6x+5

OpenStudy (cwrw238):

first add 4 to both sides of the formula y + 4 = (x - 3)^2 now take square root of both sides x-3 = +/- sqrt(y + 4) can you continue

OpenStudy (anonymous):

Nope. Im suppose to be solving system of equations.

OpenStudy (anonymous):

I dont get how you would square root (y+4)

OpenStudy (cwrw238):

sqrt(x + 4) is called a radical or in UK we say Surd - it cannot be simplified anymore

OpenStudy (cwrw238):

if you are solving a system of equations where are the others?

OpenStudy (cwrw238):

square root of ( x -3)^2 is the square of ( x - 3) so by definition its square root is z - 3

OpenStudy (anonymous):

I am suppose to be doing systems of equations. It says : On the set of axes provided,solve the following system of equations graphically for all values of X and Y. Theres two of them: y=(x-3)^2-4 and 4y+12x=20 which I already solved.

OpenStudy (anonymous):

Im just stuck on the first equation

OpenStudy (cwrw238):

right - to solve them graphically draw them om your graph paper and where they intersect are the values of X and Y

OpenStudy (anonymous):

Right, but when I show my work, I have to do x=-b/2a then do the X Y chart

OpenStudy (cwrw238):

you need to plot various values of X and Y for both equations

OpenStudy (anonymous):

Exactly but I keep getting 30's and 70's as the y value

OpenStudy (anonymous):

I get y=x^2-6X+5

OpenStudy (cwrw238):

to get values of -b/2a you need to rearrange them to the form y = ax^2 + bx + c

OpenStudy (anonymous):

Thats what Ive been trying to seek help with!

OpenStudy (cwrw238):

sorry I'd really like to help by I have to go right now

OpenStudy (anonymous):

Ok thnx

OpenStudy (anonymous):

Someone help meeeeeee pleaseeeeee

OpenStudy (radar):

Solve for x, then substitute the values you obtain for x and solve for y.

OpenStudy (anonymous):

How do I rearrange the equation?

OpenStudy (anonymous):

Its has to be in ax^2+bx+c

OpenStudy (radar):

I would start with the 2nd equation, 4y + 12x = 20 and express y in terms of x like thios 4y = -12x + 20 simply by dividing both sides by 4 y=-3x + 5 Did you follow so far?

OpenStudy (anonymous):

Yes. I got that. Thnk

OpenStudy (anonymous):

Now, when I do the other equation I get x^2-6x+5 but Im pretty sure that is wrong.

OpenStudy (radar):

Now looking at the first equation we see that y = (x-3)^2 -4 substituting or new expression for y we now have: -3x+5 = (x-3)^2 -4 do you understand that relationship?

OpenStudy (anonymous):

Hm... Never knew we had to combine them. But yes.

OpenStudy (radar):

Now go ahead and square the (x-3) and write the equations as: -3x+5 = x^2 -6x + 9 -4 combining and simplify x^2 - 9x = 0 do you want me to show how that comes about?

OpenStudy (radar):

I will continue, factor the x^2 - 9x =0 and get x(x-9) =0 solve for x getting x=0 and x=9 Theses are the two values of x.

OpenStudy (radar):

Now substitute the value of x in one of the original equations (suggest 2nd) and solve for a corresponding y value.

OpenStudy (anonymous):

Wouldn't it be x^2-3x+0?

OpenStudy (anonymous):

When you simplify, I get that because you're adding 3X to -6X

OpenStudy (radar):

Lets check: -3x + 5 = x^2 -6x + 9 - 4 combine similar terms on each side -3x + 5 = x^2 - 6x + 5 now subtract 5 and add 6x to both sides getting 3x =x^2 Now subtract x^2 from both sides -x^2 + 3x = 0 now multiply both sides by -1 (just to get a +x^2 x^2 - 3x = 0 Note I did correct an error, I was wrong about the -9.

OpenStudy (radar):

Now factor x(x-3)=0 x=0 x=3 O.K so far?

OpenStudy (anonymous):

Alrighty. I got that .

OpenStudy (radar):

Yes you found my error

OpenStudy (anonymous):

Yes

OpenStudy (radar):

Lets substitute in the 2nd orig equation and find y When x = 0 4y + 0 = 20 y=5, so one point of intersection is (0,5) When x = 3 4y + 36 = 20, 4y = -16, y=-4 the other point is (3,-4)

OpenStudy (radar):

Now the big test in order to verify those x values we substitute in the first original equation. Cross your fingers lol.

OpenStudy (anonymous):

Right. But when I do -b/2a i get a fraction.

OpenStudy (anonymous):

Then when I do the X Y chart it makes no sense

OpenStudy (radar):

y = (0-3)^2 -4 , y=9-4=5 y= (3-3)^2 - 4 y= 0^2 -4 y=0-4 y=-4 Yeaa it is correct.

OpenStudy (anonymous):

Yea i got that too

OpenStudy (radar):

We used the substitution method in solving this, we substituted for y. What method did they tell you to uses?

OpenStudy (anonymous):

It says to solve the following system of equations graphically for all values of x and y

OpenStudy (anonymous):

But I didnt understand how to rearrange the first equation in order to do -b/2a then do your tables.

OpenStudy (radar):

Oh, well you are going to get one curve and a line that will cross the curve at those two points.

OpenStudy (radar):

The first equation y = (x-3)^2 -4 will give you the curve go ahead and work that out getting: y=x^2 -6x + 9 -4 or y = x^2 -6x +5 that is a parabola

OpenStudy (anonymous):

Right.but I have to show -b/2a and the tables

OpenStudy (anonymous):

Yes thats what i got but when i plug it in the calculator and see the table the y's are high

OpenStudy (radar):

-b/2a = -3 I don't know why they are wanting you to plot that?????

OpenStudy (anonymous):

Exactly. So I thought I was doing something wrong.

OpenStudy (radar):

 NOw do the 2nd one and you will get a linear line. where it crosses the curve are the points we have already located and is the solution of the system